OpenStudy (anonymous):

How do I graph the polynomial function f(x)=x^4-4?

OpenStudy (homeworksucks):

Plot the points using a table of values?

OpenStudy (anonymous):

what do I use for x is f(x) is y?

OpenStudy (homeworksucks):

yeah

OpenStudy (homeworksucks):

f(2)=2^4-4 f(2)=16-4 f(2)=12

OpenStudy (anonymous):

Ok , Thank you!

OpenStudy (anonymous):

plotting points and using a table of values is the baby way to graph something

OpenStudy (precal):

you could take the parent function y=x^4 and move it down 4 units